During 2020, Sheridan Furniture Company purchases a carload of wicker chairs. The manufacturer sells the chairs to Sheridan for a lump sum of $83,790 because it is discontinuing manufacturing operations and wishes to dispose of its entire stock. Three types of chairs are included in the carload. The three types and the estimated selling price for each are listed below.

Type

No. of Chairs

Estimated Selling
Price Each

Lounge chairs

560 $90

Armchairs

420 80

Straight chairs

980 50


During 2020, Sheridan sells 280 lounge chairs, 140 armchairs, and 168 straight chairs.

What is the amount of gross profit realized during 2020? What is the amount of inventory of unsold straight chairs on December 31, 2020? (Round cost per chair to 2 decimal places, e.g. 78.25 and final answer to 0 decimal places, e.g. 5,845.)

Gross profit realized during 2020

$enter a dollar amount

Amount of inventory of unsold straight chairs
Answer #1
Gross profit realized during 2020 = $ 16,576
Amount of Inventory of unsold Straight chairs = $ 25,578
Workings:
Type No. Of chairs Estimated selling price Estimated Revenue Cost per Chair Chairs Actually sold Cost of goods sold Actual Revenue Profit realized Ending Units Cost of Ending Inventory
(i) (ii) (iii) = (i)*(ii) (iv) = (83790*(i))/(133000*(ii)) (v) (vi) = (iv)*(v) (vii) = (v)*(ii) (viii) = (vii) - (vi) (ix) = (i) - (v) (ix)*(iv)
Lounge Chairs 560 $          90 $       50,400 $                 56.70 280 $ 15,876 $ 25,200 $    9,324 280 $ 15,876
Arm Chairs 420 $          80 $       33,600 $                 50.40 140 $    7,056 $ 11,200 $    4,144 280 $ 14,112
Straight Chairs 980 $          50 $       49,000 $                 31.50 168 $    5,292 $    8,400 $    3,108 812 $ 25,578
$   1,33,000 $ 28,224 $ 44,800 $ 16,576